An aluminium flagpole is 33 m high.  By how much does its length increase as the temperature increases by 15 C°?

Grade:upto college level

1 Answers

Abu Nayeem
6 years ago
The coefficient of thermal expansion for aluminum, α = 22.2(10)^(-6) m/m-K. Given a change in temperature, ΔT = 15 °C = 15 K, and a flag pole length of 33 meters, the total increase in length is: 
ΔL = αLΔT = (22.2(10)^(-6) m/m-K)(33 m)(15 K) = 0.011 m = 11 mm